### v1

# Flavino Dark Matter in a Non-Abelian Discrete Flavor Model

Flavino Dark Matter
in a Non-Abelian Discrete Flavor Model
Takaaki Nomura $1$nomura@scu.edu.cn,  Yusuke Shimizu 2,3222shimizu.yusuke@kaishi-pu.ac.jp,  Towa Takahashi 2333towa@muse.sc.niigata-u.ac.jp
1College of Physics, Sichuan University, Chengdu 610065, China 2Graduate School of Science and Technology, Niigata University,
Niigata, 950-2181, Japan
3Department of Information, Kaishi Professional University,
Niigata 950-0916, Japan
( Abstract
We study a relic density of the flavino dark matter in modified Altarelli and Feruglio $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. We calculate the Lagrangian from the superpotential in the model. In estimating relic density, we consider the relevant interactions from the Lagrangian that realize the vacuum expectation value alignments and charged lepton masses where we assume that the supersymmetry breaking effects are small for flavon sector. As a result, we find the degenerate masses between the lightest “flavon” and “flavino”, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters are estimated from relic density calculation taking a constraint from lepton flavor violation into account. We also briefly discuss other dark matter physics such as direct detection, indirect detection and collider search. )
1 Introduction
The standard model (SM) is one of the successful ones with discovery of the Higgs boson. However, there are some mysterious problems in the SM, e.g. the origin of the generations for particle physics. In the SM, one cannot explain the differences of the mixing angles and masses for the quarks and leptons. Yukawa couplings are completely free parameters so that the mixing angles and masses cannot be predicted in the SM. Furthermore, the neutrinos are massless for the renormalizable operators. One of the attractive phenomena to solve the problems is the neutrino oscillation which is the evidence of the beyond the SM. Actually, the neutrino oscillation experiments provide us the important informations which are the two neutrino mass squared differences and two large mixing angles. The reactor neutrino experiments reported the last mixing angle which is non-zero value [1, 2]. Furthermore, the neutrino oscillation experiments go to new stage to measure the Dirac CP phase in the lepton sector. The neutrinos are thus windows to the new physics beyond the SM.
The flavor symmetry can be applied to the three generations of the SM fermions. The Froggatt-Nielsen mechanism [3] which was proposed by C. D. Froggatt and H. B. Nielsen applies a global $U(1)_{FN}$ symmetry. Thanks to the symmetry, one can naturally explain the fermion mass hierarchies. The non-Abelian discrete symmetry (See for the review [4]-[8].) can naturally explain the lepton mixing angles which are tri-bimaximal mixing (TBM) [9, 10] before the reactor experiments reported the non-zero reactor angle $\theta_{13}$ [1, 2]. After measurements of the $\theta_{13}$, many authors have studied the breaking or deviation from the TBM [11]-[33] or other patterns of the lepton mixing angles, e.g. tri-bimaximal-Cabibbo mixing [34, 35]. One of the successful flavor models was proposed by G. Altarelli and F. Feruglio [36, 37]. The Altarelli and Feruglio (AF) model realizes the TBM by using the non-Abelian discrete symmetry $A_{4}$. They mainly introduced two SM gauge singlet scalar fields so-called “flavons” taking the vacuum expectation value (VEV) alignments of the $A_{4}$ triplets as $(1,0,0)$ and $(1,1,1)$, which naturally explains that the charged lepton mass matrix is diagonal and neutrino mixing is TBM, respectively. However, the correct VEV alignments cannot be driven from the potential analysis. Then, they adopted the supersymmetry (SUSY) and introduced the so-called “driving” fields.
Introduction of the SUSY can solve another mysterious problem that is existence of dark matter (DM) in the universe since the lightest supersymmetric particle (LSP) is stable due to $R$-parity and it can be DM if it is neutral one [38, 39]. In the minimal supersimmetric SM (MSSM), the lightest neutralino is a good DM candidate and its properties are well discussed such as parameter space explaining observed relic density [40]. Also the neutralino DM is searched for in direct/indirect detection and collider experiments [41, 42, 43, 44, 45, 46]. However we have not seen any clear evidence of a DM in these experiments yet and nature of DM is an open question. Interestingly, in SUSY models with flavons, we have new candidate of DM that is the superpartner of the flavon which we call the ”flavino” [47]. The lightest flavino can be DM if it is the LSP since flavinos are neutral particles. Thus it is worth exploring the possibility of flavino DM.
In this paper, we discuss a relic density of the flavino DM in modified AF $A_{4}$ model [27],[48]-[51] which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In estimating relic density, we consider relevant interactions from superpotential that realize the VEV alignments and charged lepton masses where we assume SUSY breaking effects are small for flavon sector. As a result, we find degenerate mass between the lightest flavon and flavino, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters are estimated from relic density calculation taking a constraint from lepton flavor violation (LFV) into account. We also briefly discuss other DM physics such as direct detection, indirect detection and collider search.
This paper is organized as follows. In Section 2, we briefly introduce the modified AF $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. In Section 3, we discuss the relic density of the flavino DM calculating its annihilation cross section. Section 4 is devoted to a summary. We show the basis and relevant multiplication rule of the $A_{4}$ group in Appendix A. The LFV constraint for our model is shown in Appendix B.
2 Model
In this section, we briefly introduce an $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ model [27],[48]-[51]. Three generations of the left-handed lepton doublet superfields $\Phi_{\ell}=(\Phi_{\ell 1},\Phi_{\ell 2},\Phi_{\ell 3})$ are assigned to the triplet as $\bf 3$ for $A_{4}$ symmetry. The right-handed charged lepton superfields $\Phi_{e_{R}^{c}}$, $\Phi_{\mu_{R}^{c}}$, and $\Phi_{\tau_{R}^{c}}$ are assigned to the singlets as $\bf 1$, $\bf 1^{\prime\prime}$, and $\bf 1^{\prime}$ for $A_{4}$ symmetry, respectively. One of the Higgs doublet superfields is defined as $\Phi_{d}$ and assigned to the singlet as $\bf 1$ for $A_{4}$ symmetry. We introduce a superfields $\Phi_{T}=(\Phi_{T1},\Phi_{T2},\Phi_{T3})$ which are so-called “flavon superfields” and assigend to the triplet as $\bf 3$ for $A_{4}$ symmetry. In order to obtain the relevant couplings, we add the $Z_{3}$ symmetry. We also add the $U(1)_{R}$ symmetry so that we can generate the VEVs and VEV alignments through the $F$-terms by coupling flavon superfields to so-called “driving superfields” $\Phi_{0}^{T}=(\Phi_{01}^{T},\Phi_{02}^{T},\Phi_{03}^{T})$ which are assigned to triplet as $\bf 3$ for $A_{4}$ symmetry and carry the $R$ charge $+2$ under $U(1)_{R}$ symmmetry. The charge assignments of $SU(2)_{L}$, $A_{4}$, $Z_{3}$, and $U(1)_{R}$ are summarized in Tab. 1.
$\Phi_{\ell}$ $\Phi_{e_{R}^{c}}$ $\Phi_{\mu_{R}^{c}}$ $\Phi_{\tau_{R}^{c}}$ $\Phi_{d}$ $\Phi_{T}$ $\Phi_{0}^{T}$
$SU(2)_{L}$ $2$ $1$ $1$ $1$ $2$ $1$ $1$
$A_{4}$ ${\bf 3}$ ${\bf 1}$ ${\bf 1^{\prime\prime}}$ ${\bf 1^{\prime}}$ ${\bf 1}$ ${\bf 3}$ ${\bf 3}$
$Z_{3}$ $\omega$ $\omega^{2}$ $\omega^{2}$ $\omega^{2}$ $1$ $1$ $1$
$U(1)_{R}$ $1$ $1$ $1$ $1$ 00 00 $2$
Table 1: The charge assignments of $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry in our model.
In these setup, we can now write douwn the superpotential for respecting $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry at the leading order in terms of the $A_{4}$ cut-off scale $\Lambda$ as
$w_{\Phi_{T}}$ $=w_{\ell}+w_{d}^{T},$
$w_{\ell}$ $=y_{e}\Phi_{T}\Phi_{\ell}\Phi_{e_{R}^{c}}\Phi_{d}/\Lambda+y_{\mu}\Phi_{T}\Phi_{\ell}\Phi_{\mu_{R}^{c}}\Phi_{d}/\Lambda+y_{\tau}\Phi_{T}\Phi_{\ell}\Phi_{\tau_{R}^{c}}\Phi_{d}/\Lambda,$
$w_{d}^{T}$ $=-M\Phi_{0}^{T}\Phi_{T}+g\Phi_{0}^{T}\Phi_{T}\Phi_{T},$ (1)
where $\Phi_{i}$’s are chiral superfields $\Phi_{i}=\phi_{i}+\sqrt{2}\theta\psi_{i}+\theta\theta F_{i}$, $y_{j}$’s are Yukawa couplings, $M$ is mass paremeter for the flavon superfields $\Phi_{T}$ and $\Phi_{0}^{T}$  [^4] Hereafter we also call deriving field $\Phi_{0}^{T}$ as flavon field for simplicity., and $g$ is trilinear coupling for flavon superfields. The superpotential in Eq. (1) is rewritten as
$w_{\ell}$ $=y_{e}(\Phi_{T1}\Phi_{\ell_{1}}+\Phi_{T2}\Phi_{\ell_{3}}+\Phi_{T3}\Phi_{\ell_{2}})\Phi_{e_{R}^{c}}\Phi_{d}/\Lambda$
$+y_{\mu}(\Phi_{T3}\Phi_{\ell_{3}}+\Phi_{T1}\Phi_{\ell_{2}}+\Phi_{T2}\Phi_{\ell_{1}})\Phi_{\mu_{R}^{c}}\Phi_{d}/\Lambda$
$+y_{\tau}(\Phi_{T2}\Phi_{\ell_{2}}+\Phi_{T3}\Phi_{\ell_{1}}+\Phi_{T1}\Phi_{\ell_{3}})\Phi_{\tau_{R}^{c}}\Phi_{d}/\Lambda,$
$w_{d}^{T}$ $=-M(\Phi_{01}^{T}\Phi_{T1}+\Phi_{02}^{T}\Phi_{T3}+\Phi_{03}^{T}\Phi_{T2})$
$+\frac{2}{3}g\left[\Phi_{01}^{T}(\Phi_{T1}\Phi_{T1}-\Phi_{T2}\Phi_{T3})+\Phi_{02}^{T}(\Phi_{T2}\Phi_{T2}-\Phi_{T3}\Phi_{T1})+\Phi_{03}^{T}(\Phi_{T3}\Phi_{T3}-\Phi_{T1}\Phi_{T2})\right].$ (2)
Then, the Lagrangian $\mathcal{L}_{\Phi_{T}}$ for our model is given as
$\mathcal{L}_{\Phi_{T}}$ $=\int d^{2}\theta w_{\Phi_{T}}+\int d^{2}\bar{\theta}\bar{w}_{\Phi_{T}}-V_{\Phi_{T}},$
$V_{\Phi_{T}}$ $=V_{\ell}+V_{T},$ (3)
where the scalar potential $V_{\ell}$ and $V_{T}$ are obtained from $w_{d}^{T}$ and $w_{\ell}$ respectively. The scalar potential $V_{T}$ is obtained as
$V_{T}$ $=\sum_{X}\left|\frac{\partial w_{d}^{T}}{\partial X}\right|^{2}$
$=\left|-M\phi_{T1}+\frac{2}{3}g(\phi_{T1}\phi_{T1}-\phi_{T2}\phi_{T3})\right|^{2}$
$+\left|-M\phi_{T3}+\frac{2}{3}g(\phi_{T2}\phi_{T2}-\phi_{T3}\phi_{T1})\right|^{2}$
$+\left|-M\phi_{T2}+\frac{2}{3}g(\phi_{T3}\phi_{T3}-\phi_{T1}\phi_{T2})\right|^{2}$
$+\left|-M\phi_{01}^{T}+\frac{2}{3}g(2\phi_{01}^{T}\phi_{T1}-\phi_{02}^{T}\phi_{T3}-\phi_{03}^{T}\phi_{T2})\right|^{2}$
$+\left|-M\phi_{03}^{T}+\frac{2}{3}g(2\phi_{02}^{T}\phi_{T2}-\phi_{03}^{T}\phi_{T1}-\phi_{01}^{T}\phi_{T3})\right|^{2}$
$+\left|-M\phi_{02}^{T}+\frac{2}{3}g(2\phi_{03}^{T}\phi_{T3}-\phi_{01}^{T}\phi_{T2}-\phi_{02}^{T}\phi_{T1})\right|^{2},$ (4)
where $X=\phi_{T1}$, $\phi_{T2}$, $\phi_{T3}$, $\phi_{01}^{T}$, $\phi_{02}^{T}$, and $\phi_{03}^{T}$. On the other hand, the scalar potential $V_{\ell}$ is written by
$V_{\ell}$ $=\sum_{Y}\left|\frac{\partial w_{\ell}}{\partial Y}\right|^{2}$
$=\left|(y_{e}\tilde{\phi}_{\ell_{1}}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\tilde{\phi}_{\ell_{2}}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\tilde{\phi}_{\ell_{3}}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|(y_{e}\tilde{\phi}_{\ell_{3}}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\tilde{\phi}_{\ell_{1}}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\tilde{\phi}_{\ell_{2}}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|(y_{e}\tilde{\phi}_{\ell_{2}}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\tilde{\phi}_{\ell_{3}}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\tilde{\phi}_{\ell_{1}}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|(y_{e}\phi_{T1}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\phi_{T2}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\phi_{T3}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|(y_{e}\phi_{T3}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\phi_{T1}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\phi_{T2}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|(y_{e}\phi_{T2}\tilde{\phi}_{e_{R}^{c}}+y_{\mu}\phi_{T3}\tilde{\phi}_{\mu_{R}^{c}}+y_{\tau}\phi_{T1}\tilde{\phi}_{\tau_{R}^{c}})h_{d}/\Lambda\right|^{2}$
$+\left|y_{e}(\phi_{T1}\tilde{\phi}_{\ell_{1}}+\phi_{T2}\tilde{\phi}_{\ell_{3}}+\phi_{T3}\tilde{\phi}_{\ell_{2}})h_{d}/\Lambda\right|^{2}$
$+\left|y_{\mu}(\phi_{T3}\tilde{\phi}_{\ell_{3}}+\phi_{T1}\tilde{\phi}_{\ell_{2}}+\phi_{T2}\tilde{\phi}_{\ell_{1}})h_{d}/\Lambda\right|^{2}$
$+\left|y_{\tau}(\phi_{T2}\tilde{\phi}_{\ell_{2}}+\phi_{T3}\tilde{\phi}_{\ell_{1}}+\phi_{T1}\tilde{\phi}_{\ell_{3}})h_{d}/\Lambda\right|^{2}$
$+\left|\left[y_{e}(\phi_{T1}\tilde{\phi}_{\ell_{1}}+\phi_{T2}\tilde{\phi}_{\ell_{3}}+\phi_{T3}\tilde{\phi}_{\ell_{2}})\tilde{\phi}_{e_{R}^{c}}+y_{\mu}(\phi_{T3}\tilde{\phi}_{\ell_{3}}+\phi_{T1}\tilde{\phi}_{\ell_{2}}+\phi_{T2}\tilde{\phi}_{\ell_{1}})\tilde{\phi}_{\mu_{R}^{c}}\right.\right.$
$+\left.\left.y_{\tau}(\phi_{T2}\tilde{\phi}_{\ell_{2}}+\phi_{T3}\tilde{\phi}_{\ell_{1}}+\phi_{T1}\tilde{\phi}_{\ell_{3}})\tilde{\phi}_{\tau_{R}^{c}}\right]/\Lambda\right|^{2},$ (5)
where $Y=\phi_{T1}$, $\phi_{T2}$, $\phi_{T3}$, $\tilde{\phi}_{\ell_{1}}$, $\tilde{\phi}_{\ell_{2}}$, $\tilde{\phi}_{\ell_{3}}$, $\tilde{\phi}_{e_{R}^{c}}$, $\tilde{\phi}_{\mu_{R}^{c}}$, $\tilde{\phi}_{\tau_{R}^{c}}$, and $h_{d}$ with $\tilde{\phi}$ scalar fields being slepton ones. The F-term contribution for our model is calculated as
$\int d^{2}\theta w_{\Phi_{T}}$ $=y_{e}(\phi_{T1}\ell_{1}+\phi_{T2}\ell_{3}+\phi_{T3}\ell_{2})e_{R}^{c}h_{d}/\Lambda$
$+y_{\mu}(\phi_{T3}\ell_{3}+\phi_{T1}\ell_{2}+\phi_{T2}\ell_{1})\mu_{R}^{c}h_{d}/\Lambda$
$+y_{\tau}(\phi_{T2}\ell_{2}+\phi_{T3}\ell_{1}+\phi_{T1}\ell_{3})\tau_{R}^{c}h_{d}/\Lambda$
$+y_{e}(\tilde{\psi}_{\phi_{T1}}\tilde{\phi}_{\ell_{1}}+\tilde{\psi}_{\phi_{T2}}\tilde{\phi}_{\ell_{3}}+\tilde{\psi}_{\phi_{T3}}\tilde{\phi}_{\ell_{2}})e_{R}^{c}h_{d}/\Lambda$
$+y_{\mu}(\tilde{\psi}_{\phi_{T3}}\tilde{\phi}_{\ell_{3}}+\tilde{\psi}_{\phi_{T1}}\tilde{\phi}_{\ell_{2}}+\tilde{\psi}_{\phi_{T2}}\tilde{\phi}_{\ell_{1}})\mu_{R}^{c}h_{d}/\Lambda$
$+y_{\tau}(\tilde{\psi}_{\phi_{T2}}\tilde{\phi}_{\ell_{2}}+\tilde{\psi}_{\phi_{T3}}\tilde{\phi}_{\ell_{1}}+\tilde{\psi}_{\phi_{T1}}\tilde{\phi}_{\ell_{3}})\tau_{R}^{c}h_{d}/\Lambda$
$+y_{e}(\tilde{\psi}_{\phi_{T1}}\ell_{1}+\psi_{\phi_{T2}}\ell_{3}+\psi_{\phi_{T3}}\ell_{2})\tilde{\phi}_{e_{R}^{c}}h_{d}/\Lambda$
$+y_{\mu}(\tilde{\psi}_{\phi_{T3}}\ell_{3}+\psi_{\phi_{T1}}\ell_{2}+\psi_{\phi_{T2}}\ell_{1})\tilde{\phi}_{\mu_{R}^{c}}h_{d}/\Lambda$
$+y_{\tau}(\tilde{\psi}_{\phi_{T2}}\ell_{2}+\psi_{\phi_{T3}}\ell_{1}+\psi_{\phi_{T1}}\ell_{3})\tilde{\phi}_{\tau_{R}^{c}}h_{d}/\Lambda$
$+y_{e}(\tilde{\psi}_{\phi_{T}1}\tilde{\phi}_{\ell_{1}}+\tilde{\psi}_{\phi_{T}2}\tilde{\phi}_{\ell_{3}}+\tilde{\psi}_{\phi_{T}3}\tilde{\phi}_{\ell_{2}})\tilde{\phi}_{e_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\mu}(\tilde{\psi}_{\phi_{T}3}\tilde{\phi}_{\ell_{3}}+\tilde{\psi}_{\phi_{T}1}\tilde{\phi}_{\ell_{2}}+\tilde{\psi}_{\phi_{T}2}\tilde{\phi}_{\ell_{1}})\tilde{\phi}_{\mu_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\tau}(\tilde{\psi}_{\phi_{T}2}\tilde{\phi}_{\ell_{2}}+\tilde{\psi}_{\phi_{T}3}\tilde{\phi}_{\ell_{1}}+\tilde{\psi}_{\phi_{T}1}\tilde{\phi}_{\ell_{3}})\tilde{\phi}_{\tau_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{e}(\phi_{T1}\ell_{1}+\phi_{T2}\ell_{3}+\phi_{T3}\ell_{2})\tilde{\phi}_{e_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\mu}(\phi_{T3}\ell_{3}+\phi_{T1}\ell_{2}+\phi_{T2}\ell_{1})\tilde{\phi}_{\mu_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\tau}(\phi_{T2}\ell_{2}+\phi_{T3}\ell_{1}+\phi_{T1}\ell_{3})\tilde{\phi}_{\tau_{R}^{c}}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{e}(\phi_{T1}\tilde{\phi}_{\ell_{1}}+\phi_{T2}\tilde{\phi}_{\ell_{3}}+\phi_{T3}\tilde{\phi}_{\ell_{2}})e_{R}^{c}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\mu}(\phi_{T3}\tilde{\phi}_{\ell_{3}}+\phi_{T1}\tilde{\phi}_{\ell_{2}}+\phi_{T2}\tilde{\phi}_{\ell_{1}})\mu_{R}^{c}\tilde{\psi}_{h_{d}}/\Lambda$
$+y_{\tau}(\phi_{T2}\tilde{\phi}_{\ell_{2}}+\phi_{T3}\tilde{\phi}_{\ell_{1}}+\phi_{T1}\tilde{\phi}_{\ell_{3}})\tau_{R}^{c}\tilde{\psi}_{h_{d}}/\Lambda$
$-M(\tilde{\psi}_{\phi_{01}^{T}}\tilde{\psi}_{\phi_{T1}}+\tilde{\psi}_{\phi_{02}^{T}}\tilde{\psi}_{\phi_{T3}}+\tilde{\psi}_{\phi_{03}^{T}}\tilde{\psi}_{\phi_{T2}})$
$+\frac{2}{3}g\left[\phi_{01}^{T}(\tilde{\psi}_{\phi_{T1}}\tilde{\psi}_{\phi_{T1}}-\tilde{\psi}_{\Phi_{T2}}\tilde{\psi}_{\phi_{T3}})+\phi_{02}^{T}(\tilde{\psi}_{\phi_{T2}}\tilde{\psi}_{\phi_{T2}}-\tilde{\psi}_{\phi_{T3}}\tilde{\psi}_{\phi_{T1}})\right.$
$+\phi_{03}^{T}(\tilde{\psi}_{\phi_{T3}}\tilde{\psi}_{\phi_{T3}}-\tilde{\psi}_{\phi_{T1}}\tilde{\psi}_{\phi_{T2}})$
$+\tilde{\psi}_{\phi_{01}^{T}}(\phi_{T1}\tilde{\psi}_{\phi_{T1}}-\phi_{T2}\tilde{\psi}_{\phi_{T3}})+\tilde{\psi}_{\phi_{02}^{T}}(\phi_{T2}\tilde{\psi}_{\phi_{T2}}-\phi_{T3}\tilde{\psi}_{\phi_{T1}})+\tilde{\psi}_{\phi_{03}^{T}}(\phi_{T3}\tilde{\psi}_{\phi_{T3}}-\phi_{T1}\tilde{\psi}_{\phi_{T2}})$
$+\left.\tilde{\psi}_{\phi_{01}^{T}}(\tilde{\psi}_{\phi_{T1}}\phi_{T1}-\tilde{\psi}_{\phi_{T2}}\phi_{T3})+\tilde{\psi}_{\phi_{02}^{T}}(\tilde{\psi}_{\phi_{T2}}\phi_{T2}-\tilde{\psi}_{\phi_{T3}}\phi_{T1})+\tilde{\psi}_{\phi_{03}^{T}}(\tilde{\psi}_{\phi_{T3}}\phi_{T3}-\tilde{\psi}_{\phi_{T1}}\phi_{T2})\right],$ (6)
where $\tilde{\psi}$ fields correspond to flavinos. The VEV alignments of $\phi_{T}$ and $\phi_{0}^{T}$ are obtained from the potential minimum condition, $V_{T}=0$, as in Ref. [50];
$\langle\phi_{T}\rangle=v_{T}(1,0,0),\qquad v_{T}=\frac{3M}{2g},\qquad\langle\phi_{0}^{T}\rangle=(0,0,0).$ (7)
The charged lepton masses are then obtained after $h_{d}$ developing its VEV, $\langle h_{d}\rangle=v_{d}$, as $m_{\{e,\mu,\tau\}}=y_{\{e,\mu,\tau\}}v_{d}v_{T}/\Lambda$. After taking the VEV alignments, e.g. the $A_{4}$ triplet flavon ($\phi_{T1}$, $\phi_{T2}$, $\phi_{T3}$)=($v_{T}+\varphi_{T1}$, $\phi_{T2}$, $\phi_{T3}$), we find the mass terms of flavons and flavinos as follows,
$\mathcal{L}_{\Phi_{T}}\supset$ $M^{2}|\varphi_{T1}|^{2}+4M^{2}|\phi_{T2}|^{2}+4M^{2}|\phi_{T3}|^{2}+M^{2}|\phi_{01}^{T}|^{2}+4M^{2}|\phi_{02}^{T}|^{2}+4M^{2}|\phi_{03}^{T}|^{2}$
$-M\tilde{\psi}_{\phi^{T}_{01}}\tilde{\psi}_{\phi_{T1}}-2M(\tilde{\psi}_{\phi_{02}^{T}}\tilde{\psi}_{\phi_{T3}}+\tilde{\psi}_{\phi_{03}^{T}}\tilde{\psi}_{\phi_{T2}})+h.c.\,.$ (8)
Notice that pairs of chiral fermion components $\{\tilde{\psi}_{\phi^{T}_{01}},\tilde{\psi}_{\phi_{T1}}\}$, $\{\tilde{\psi}_{\phi_{02}^{T}},\tilde{\psi}_{\phi_{T3}}\}$ and $\{\tilde{\psi}_{\phi_{03}^{T}},\tilde{\psi}_{\phi_{T2}}\}$ construct Dirac mass terms. We also find that the mass ratio of flavons(flavinos) is $1:2:2$, and flavino and flavon masses are degenerate without considering SUSY breaking effects. In this work the DM candidate is the lightest flavino which is Dirac fermion with mass $M$. We rewrite it as
$X_{R}\equiv\tilde{\psi}^{c}_{\phi^{T}_{01}},\qquad X_{L}\equiv\tilde{\psi}_{\phi_{T1}},$ (9)
where "$c"$ indicates charge conjugation and $X$ is identified as Dirac fermion. Here we focus on interactions among the DM $X$ and the lightest flavons $\varphi_{T1}$ and $\phi^{T}_{01}$ since it provide dominant contributions in relic density estimation as we discuss in next section. The relevant interactions are summarized by
$\mathcal{L}_{\Phi_{T}}\supset$ $\frac{M}{v_{T}}\left[2\varphi_{T1}\overline{X_{R}}X_{L}+\phi_{01}^{T}\overline{X^{c}_{L}}X_{L}+h.c.\right]$
$\frac{M^{2}}{v_{T}}\left[\varphi_{T1}\varphi_{T1}^{*}\varphi_{T1}^{*}+c.c.\right]-\frac{2M^{2}}{v_{T}}\left[\phi_{01}^{T}\phi_{01}^{T*}\varphi_{T1}^{*}+c.c.\right].$ (10)
These interactions induce DM annihilation processes that determine relic density.
3 Dark matter physics
In this section we discuss relic density of flavino DM calculating its annihilation cross section. Other DM phenomenologies such as direct/indirect detections and collider searches are also discussed briefly.
3.1 DM annihilation cross sections
In our scenario flavino mainly annihilate into the lightest flavons, $\{\varphi_{T},\phi_{01}^{T}\}$, due to small couplings among flavino and the SM leptons. There are many possible processes of such annihilation processes as
$\overline{X}X\rightarrow\{\varphi_{T1}\varphi_{T1},\overline{\varphi_{T1}}\ \overline{\varphi_{T1}},\overline{\phi^{T}_{01}}\phi^{T}_{01},\overline{\varphi_{T1}}\varphi_{T1}\},$
$XX\rightarrow\overline{\phi^{T}_{01}}\ \overline{\varphi_{T1}},\quad\overline{X}\ \overline{X}\rightarrow\phi^{T}_{01}\varphi_{T1}.$ (11)
Although many diagrams can be drawn from the Lagrangian, we find the dominant contributions are $\overline{X}X\rightarrow\{\varphi_{T1}\varphi_{T1},\overline{\varphi_{T1}}\ \overline{\varphi_{T1}}\}$ and $XX\rightarrow\overline{\phi^{T}_{01}}\ \overline{\varphi_{T1}}(\overline{X}\ \overline{X}\rightarrow\phi^{T}_{01}\varphi_{T_{1}})$ where the other processes are suppressed by $p^{2}/M^{2}$ factor with $p$ being magnitude of DM momentum. In the following we summarize the cross sections of the dominant processes.
(i) $X\bar{X}\longrightarrow\varphi_{T1}\varphi_{T1}$:
In this process, all channels are present and relevant matrix elements are
(a) s-channel
(b) t-channel
(c) t-channel
Figure 1: Diagrams for the process $\bar{X}X\to\varphi_{T1}\varphi_{T1}$.
$\mathcal{M}_{s}$ $=-\left(\frac{2M}{v_{T}}\right)\left(\frac{2M^{2}}{v_{T}}\right)\frac{1}{s-M^{2}}\bar{v}_{(p_{2})}P_{R}u_{(p_{1})}$
$\mathcal{M}_{t}$ $\thicksim{\left(\frac{M}{v_{T}}\right)}^{2}\frac{1}{M}\bar{v}_{(p_{2})}P_{R}u_{(p_{1})}$
$\mathcal{M}_{u}$ $\thicksim{\left(\frac{M}{v_{T}}\right)}^{2}\frac{1}{M}\bar{v}_{(p_{2})}P_{R}u_{(p_{1})},$ (12)
where $s=(p_{1}+p_{2})^{2}$ with $p_{1(2)}$ being flavino $X(\bar{X})$ momentum in the initial state. Here, we ignored other Mandelstam variables assuming $t,u\ll M^{2}$. Since we consider non-retivistic limit, we approximate $s\thicksim 4M^{2}$. Thus, the matrix element of the process is
$\mathcal{M}\thicksim\frac{20M}{3{v_{T}}^{2}}\bar{v}_{(p_{2})}P_{R}u_{(p_{1})}.$
The cross section is
$\sigma_{X\bar{X}\rightarrow\varphi_{T1}\varphi_{T1}}\thicksim\frac{25}{72\pi}\frac{M^{2}}{{v_{T}}^{4}}\frac{s-2M^{2}}{s}.$ (13)
(ii)$X\bar{X}\longrightarrow\overline{\varphi_{T1}}\,\overline{\varphi_{T1}}$:
This process are just conjugate of $X\bar{X}\longrightarrow\varphi_{T1}\varphi_{T1}$. Thus, cross section is
$\sigma_{X\bar{X}\rightarrow\overline{\varphi_{T1}}\,\overline{\varphi_{T1}}}$ $=\sigma_{X\bar{X}\rightarrow\varphi_{T1}\varphi_{T1}}$
$\thicksim\frac{25}{72\pi}\frac{M^{2}}{{v_{T}}^{4}}\frac{s-2M^{2}}{s}.$ (14)
(iii)$XX\longrightarrow\overline{\phi^{T}_{01}}\,\overline{{\phi}_{T1}}$:
In this prosess, all channels are present. Matrix elements are
(a) s-channel
(b) t-channel
(c) t-channel
Figure 2: Diagrams for the process $XX\to\overline{\varphi_{T1}}\overline{\phi^{T}_{01}}$.
$\mathcal{M}_{s}$ $=\left(\frac{M}{v_{T}}\right)\left(\frac{2M^{2}}{v_{T}}\right)\frac{1}{s-M^{2}}\bar{v}_{(p_{2})}P_{L}u_{(p_{1})}$
$\mathcal{M}_{t}$ $\thicksim\mathcal{M}_{u}\thicksim\left(\frac{2M}{{v_{T}}^{2}}\right)\bar{v}_{(p_{2})}P_{L}u_{(p_{1})}$ (15)
This process’s matrix element is
$\mathcal{M}\thicksim\left(\frac{14M}{{3v_{T}}^{2}}\right)\bar{v}_{(p_{2})}P_{L}u_{(p_{1})}.$
The cross section is
$\sigma_{XX\rightarrow\overline{\phi^{T}_{01}}\overline{\phi_{T1}}}\thicksim\left(\frac{49}{144\pi s}\right)\left(\frac{{M}^{2}}{{v_{T}}^{4}}\right)\left(s-2M^{2}\right).$ (16)
(iv)$\bar{X}\bar{X}\longrightarrow{\phi}^{T}_{01}{\varphi}_{T1}$:
This process are just conjugate of $XX\longrightarrow\overline{\phi^{T}_{01}}\overline{\varphi_{T1}}$. Thus, cross section is
$\sigma_{\bar{X}\bar{X}\rightarrow{\phi}^{T}_{01}{\varphi}_{T1}}$ $=\sigma_{XX\rightarrow\overline{\varphi^{T}_{01}}\overline{\varphi_{T1}}}$
$\thicksim\left(\frac{49}{144\pi s}\right)\left(\frac{{M}^{2}}{{v_{T}}^{4}}\right)\left(s-2M^{2}\right).$ (17)
We apply these cross sections in DM relic density calculation.
3.2 Relic density
Relic density of flavino DM is obtained by solving Boltzmann equation for number density $n_{X}$ of DM $X$,
$\dot{n}_{X}+3Hn_{X}=\langle\sigma v\rangle(n_{X_{\rm eq}}^{2}-n_{X}^{2}),$ (18)
where $H$ is Hubble parameter, $\langle\sigma v\rangle$ is thermal average of DM annihilation cross section and $n_{X_{eq}}$ is density of $X$ in equilibrium. The annihilation cross section is the sum of cross sections discussed in the previous subsection. Relic density $\Omega_{X}h^{2}$ can be approximately estimated as [52, 53]
$\Omega_{X}h^{2}\simeq\frac{1.07\times 10^{9}}{\sqrt{g^{*}(x_{f})}M_{Pl}J(x_{f})\ {\rm[GeV]}},$ (19)
[…]
where $K_{1,2}$ denote the modified Bessel functions of the second kind of order 1 and 2.
Figure 3: Red points shows the parameters that provide observed relic density $\Omega_{X}h^{2}\simeq 0.12$. The gray region is disfavored since coupling $g$ is beyond the perturbative limit. Relic density is over abundant in the light gray region. The light blue region is excluded by the constraint from flavor violating $\tau$ decay.
In estimating relic density we apply micrOMEGAs code [54] implementing relevant interactions in the model; we compared the numerical results with analytical calculation with Eq. (19) and confirmed consistency. In Fig. 3, we show parameters that provide observed amount of relic density $\Omega_{X}h^{2}\simeq 0.12$ [55] by red points where relic density is over abundant in the light gray region and smaller than observed amount below red points. Also the gray region is disfavored since coupling $g$ becomes non-perturbative value. In addition, the light blue region is excluded by the constraint from flavor violating $\tau$ decay [48] (see Appendix B). We thus find that the range of the lightest flavino(flavon) mass $[6,6\times 10^{4}]$ [GeV] can accommodate observed relic density. Also allowed region of $v_{T}$ is roughly 30 [GeV] to $2\times 10^{4}$ [GeV].
3.3 Comments on ohter flavino DM phyiscs
Here we provide brief discussion regarding some phenomenology of flavino DM.
Direct detection:
The interactions from superpotential do not contribute to flavino-nucleon scattering at tree and one-loop level since they only have coupling terms among flavino, flavon, Higgs(Higgsinos) and leptons(sleptons); there is no flavino-flavino-Higgs interaction. Thus these interactions are not constrained by the direct detection experiments; flavino can also interact with electron via flavon $\varphi_{T1}$ exchange but cross section is tiny since flavon-electron coupling is proportional to electron mass. In fact flavino-nucleon scattering is possible via Higgs portal interactions when flavon and Higgs bosons mix through SUSY-breaking terms. In this work we assume effect of SUSY breaking is small for flavon sector to avoid current experimental bounds, e.g. XENON1T [56], PandaX-4T [57] and LUX-ZEPLIN [58].
Indirect detection:
[…]
Collider search:
Flavino DM can be searched for at collider experiments such as the LHC as the preferred scale of flavino mass is $\mathcal{O}$(10) GeV to $\mathcal{O}$(10) TeV. One possible process is slepton pair production followed by slepton decay ~$\tilde{\ell}\to\ell X$ when a slepton is the next to lightest SUSY particle. The signal of this case is the same as slepton decaying into neutralino DM [40] and it is difficult to distinguish. We can expect more specific signals of the model when heavier flavinos and flavons are lighter than sleptons. In such a case we would have cascade decay of slepton, e.g. ~$\tilde{\ell}\to\ell\overline{\tilde{\psi}_{T2}}\to\ell\overline{X}\phi_{T3}(\to\bar{\ell}^{\prime}\ell^{\prime})$, inducing multi-leptons with missing-energy signal. For these signals we may be able to reconstruct flavon mass and it helps us to confirm the model. Detailed analysis of collider signals is beyond the scope of this paper and we left it in future work.
4 Summary
We have studied the relic density of the flavino DM in modified AF $A_{4}$ model which is respecting the $SU(2)_{L}\times A_{4}\times Z_{3}\times U(1)_{R}$ symmetry. First of all, we have calculated the Lagrangian from the superpotential in the modified AF $A_{4}$ model. In estimating relic density, we have considered the relevant interactions from the Lagrangian that realize the VEV alignments and charged lepton masses where we assume SUSY breaking effects are small for flavon sector. As a result, we have found the degenerate masses between the lightest flavon and flavino, and only two parameters in the potential determines the relic density. Then allowed parameter space of these parameters have been estimated from relic density calculation taking a constraint from lepton flavor violation (LFV) into account. We thus have found that the range of lightest flavino(flavon) mass $[6,6\times 10^{4}]$ [GeV] could accommodate observed relic density. Also allowed region of $v_{T}$ has been roughly 30 [GeV] to $2\times 10^{4}$ [GeV]. We also briefly have discussed other DM physics such as direct detection, indirect detection and collider search.
Acknowledgement
We thank T. Toma for useful discussions. The work of T. N. was supported by the Fundamental Research Funds for the Central Universities. The work of Y. S. was partially supported by JSPS KAKENHI Grant No. 20H01898.
Appendix
Appendix A Multiplication rule of $A_{4}$ group
We use the multiplication rule of the $A_{4}$ triplet as follows:
$\begin{pmatrix}a_{1}\\ a_{2}\\ a_{3}\end{pmatrix}_{\bf 3}\otimes\begin{pmatrix}b_{1}\\ b_{2}\\ b_{3}\end{pmatrix}_{\bf 3}$ $=\left(a_{1}b_{1}+a_{2}b_{3}+a_{3}b_{2}\right)_{\bf 1}\oplus\left(a_{3}b_{3}+a_{1}b_{2}+a_{2}b_{1}\right)_{{\bf 1}^{\prime}}$
$\oplus\left(a_{2}b_{2}+a_{1}b_{3}+a_{3}b_{1}\right)_{{\bf 1}^{\prime\prime}}$
$\oplus\frac{1}{3}\begin{pmatrix}2a_{1}b_{1}-a_{2}b_{3}-a_{3}b_{2}\\ 2a_{3}b_{3}-a_{1}b_{2}-a_{2}b_{1}\\ 2a_{2}b_{2}-a_{3}b_{1}-a_{1}b_{3}\end{pmatrix}_{{\bf 3}}\oplus\frac{1}{2}\begin{pmatrix}a_{2}b_{3}-a_{3}b_{2}\\ a_{1}b_{2}-a_{2}b_{1}\\ a_{3}b_{1}-a_{1}b_{3}\end{pmatrix}_{{\bf 3}\ }\ .$ (21)
More details are shown in the review [5, 6].
Appendix B LFV constraint in the model
In the model some LFV charged lepton decay processes are forbidden by residual $Z_{3}$ symmetry [48]. In particular LFV muon decays are forbidden and the strongest constraint is obtained from $\tau\to\mu\mu\bar{e}$ process. The branching ratio of the process is given by
${\rm BR}(\tau\to\mu\mu\bar{e})$ $=\tau_{\tau}\frac{m_{\tau}^{5}}{3027\pi^{3}}\left(\left|\frac{m_{\tau}m_{\mu}}{v_{T}^{2}m^{2}_{\phi_{T2}}}\right|^{2}+\left|\frac{m_{\mu}m_{e}}{v_{T}^{2}m^{2}_{\phi_{T3}}}\right|^{2}\right)$
$\simeq$ $\frac{2.9\times 10^{6}\,{\rm GeV}^{8}}{v_{T}^{4}(2M)^{4}}$ (22)
where $\tau_{\tau}$ is the lifetime of tau lepton. It is compared with the current constraint by measurement at Belle experiment BR$(\tau\to\mu\mu\bar{e})<1.7\times 10^{-8}$ [61] to constrain $\{M,v_{T}\}$ parameter space.